# EgoEvGesture: Gesture Recognition Based on Egocentric Event Camera

### v1

@@ 3.3 Dataset Composition and Statistics
Figure 4: Action duration distributions across subjects under normal light (left) and low light (right) conditions. Each vertical bar represents cumulative duration for a specific action, with colored segments indicating subject-wise contributions.
(a) Average action event rates under normal light and low light conditions across 10 subjects. Each subplot represents data from a single subject, with 38 discrete actions uniformly distributed around the 360° axis. Higher variance in complex gestures indicates movement pattern diversity.
@@
(b) Subject-wise event rates under mixed, normal light and low light conditions showing individual adaptation variance.
(c) Bimanual vs unimanual event rates. Higher event density in two-hand gestures confirms asymmetric motion dynamics.
Figure 5: Comprehensive event rate statistics: (a) Action-class variability; (b) Subject-specific patterns across lighting conditions; (c) Handedness effects.
@@
In summary, the EgoEvGesture dataset offers a rich and diverse collection of egocentric gesture data, captured under varying lighting conditions, making it a valuable resource for advancing research in gesture recognition using event cameras.
4 Methodology
4.1 Architecture Overview: Dual-Motion Disentanglement Framework
@@
To address the critical challenge of ego-hand motion entanglement in event streams while capturing holistic temporal patterns, we propose a temporal pyramid architecture with three progressive processing stages (Fig. 6):
@@
Figure 6: Three-stage processing pipeline: (1) LNES converts raw events to normalized surfaces preserving recent temporal signatures (2) Blaze & VMamba blocks extract multi-scale features (3) BTSM aligns temporal context
@@
We first convert raw events into Locally Normalized Event Surfaces (LNES) that focus on the most recent temporal signature to suppress historical ego-motion artifacts while maintaining real-time performance. We then adopt Blaze & VMamba blocks to extract spatiotemporal features. Our Blaze encoder-decoder, adapted from [14] with proven efficiency in human-centric perception [16], processes each temporal bin. The VMamba block addresses sparse event processing and long-term temporal modeling through discrete state space equations (detailed in Sec. 4.2). Finally, we employ the Bidirectional temporal shift module (BTSM) to resolve temporal semantic fragmentation caused by segmented processing and perform temporal alignment (detailed in Sec. 4.3).
4.2 VMamba Block
@@
Despite the ability to adapt to high dynamic ranges, event-based gesture recognition faces challenges due to the sparsity of events, which makes it difficult for traditional methods to effectively focus on key areas. Meanwhile, the very high event resolution, while enhancing the ability to capture fine details, also makes long-term temporal understanding more challenging. To address the aforementioned challenges, we introduce the VMamba block, which addresses this modality discrepancy through adaptive temporal continuity - maintaining feature dimensions (for compatibility) while implementing dynamic state transitions. Our design philosophy centers on selective computation focusing - preserving computational resources for active temporal regions without sacrificing temporal granularity. The VMamba block operates as follows:
[…]
where: - $\Delta_{t}$ is computed via a time-aware gating mechanism, adaptively controlling temporal dependencies. - $\mathbf{B}_{t}$ and $\mathbf{C}_{t}$ are obtained through input/output projection, modulating input sparsity and serving as linear transformations. - $\mathbf{\tilde{A}}_{t}$ bridges continuous-time dynamics and discrete-time updates. - $\mathbf{\tilde{B}}_{t}$ amplifies or reduces input features based on temporal relevance. - $\mathbf{h}_{t}$ updates the hidden state through selective memory, combining temporal state propagation and sparsity-aware input integration. - $\mathbf{y}_{t}$ combines transformed hidden states with residual input connections for stable training.
@@
4.3 Bins-Temporal Shift Module (BTSM)
@@
4.3.1 Intra-Bin Fusion
The intra-bin fusion stage processes temporal relationships within each segment via channel-wise shifting, where $\frac{F}{4}$ channels are shifted left/right along temporal dimension within bins, preserving spatial locality. Given input features $X\in\mathbb{R}^{Bt\times T\times Bn\times F}$ where $Bt$=batch size, $T$=number of bins into which one action segment is divided, $Bn$=number of frames in a bin, and $F$=features, the module operates as follows:
$X_{intra}=\mathcal{T}_{local}(X)=\begin{cases}X^{left}=\text{roll}(X_{:,:,:,0:\frac{F}{4}},\text{shift}=-1,\text{dim}=2),\\ X^{right}=\text{roll}(X_{:,:,:,\frac{F}{4}:\frac{F}{2}},\text{shift}=+1,\text{dim}=2),\\ X^{static}=X_{:,:,:,\frac{F}{2}:},\\ \end{cases}$ (2)
$X_{intra}=\text{concat}(X^{left},\ X^{right},\ X^{static})\in\mathbb{R}^{Bt\times T\times Bn\times F},$ (3)
where:
  • $\text{roll}(X,\text{shift},\text{dim})$: A function that shifts the elements of $X$ along the specified dimension dim by the amount shift. A negative shift moves elements to the left, while a positive shift moves elements to the right.
  • $X^{left}$: The left-shifted features, obtained by shifting the first $\frac{F}{4}$ channels to the left along the temporal dimension.
  • $X^{right}$: The right-shifted features, obtained by shifting the next $\frac{F}{4}$ channels to the right along the temporal dimension.
  • $X^{static}$: The remaining $\frac{F}{2}$ channels that are not shifted, representing static features.
  • concat: A function that concatenates the shifted and static features along the feature dimension.
4.3.2 Inter-Bin Fusion
The Inter-Bin Fusion stage models cross-segment dependencies through bin-axis shifting. This operation implements cross-bin feature propagation while maintaining temporal ordering within bins, enabling the modeling of long-range temporal dependencies. The module operates as follows:
$\tilde{X}=X_{intra}\in\mathbb{R}^{Bt\times T\times Bn\times F},$ (4)
$X_{inter}=\mathcal{T}_{inter}(\tilde{X})=\begin{cases}\tilde{X}^{left}=\text{roll}(\tilde{X}_{:,:,:,0:\frac{F}{4}},\text{shift}=-1,\text{dim}=1),\\ \tilde{X}^{right}=\text{roll}(\tilde{X}_{:,:,:,\frac{F}{4}:\frac{F}{2}},\text{shift}=+1,\text{dim}=1),\\ \tilde{X}^{static}=\tilde{X}_{:,:,:,\frac{F}{2}:},\\ \end{cases}$ (5)
$X_{final}=\text{concat}(\tilde{X}^{left},\ \tilde{X}^{right},\ \tilde{X}^{static})\in\mathbb{R}^{Bt\times T\times Bn\times F},$ (6)
where:
  • $\tilde{X}$: The input features to the Inter-Bin Fusion stage, which are the output features from the Intra-Bin Fusion stage. It has the same dimension as $X_{intra}$, i.e., $\mathbb{R}^{Bt\times T\times Bn\times F}$.
  • $X_{inter}$: The intermediate features after applying the bin-axis shifting operation. It is a tuple containing the left-shifted, right-shifted, and static features along the bin axis.
  • $\tilde{X}^{left}$: The left-shifted features along the bin axis, obtained by shifting the first $\frac{F}{4}$ channels to the left along the bin dimension (dim=1).
  • $\tilde{X}^{right}$: The right-shifted features along the bin axis, obtained by shifting the next $\frac{F}{4}$ channels to the right along the bin dimension (dim=1).
  • $\tilde{X}^{static}$: The remaining $\frac{F}{2}$ channels that are not shifted along the bin axis, representing static features.
  • $X_{final}$: The final output features after concatenating the shifted and static features along the feature dimension. It has the same dimension as the input features $\tilde{X}$, i.e., $\mathbb{R}^{Bt\times T\times Bn\times F}$.
  • $\text{roll}(X,\text{shift},\text{dim})$ and concat: Same as in the Intra-Bin Fusion stage.
4.3.3 Mechanism Analysis
@@ 4.4 Loss Function
5 Experiments
5.1 Dataset Configuration and Evaluation Protocol
@@
5.1.1 Illumination Robustness Validation Split
To evaluate the robustness of our method under different illumination, we consider three testing scenarios for each of the 3 splits:
  • Normal light condition only: This scenario tests the model’s performance under normal light conditions, serving as a baseline to compare with other scenarios.
  • Low light condition only: Here, we assess how the model handles low light conditions, which are more challenging due to reduced visibility and altered appearance of pedestrians.
  • Mixed condition: This scenario combines both normal and low light conditions, simulating real-world situations where the model may encounter a variety of lighting conditions within the same test set.
5.1.2 Two Different Training-Test Split Methods
We implement two evaluation protocols: heterogeneous split (heter) and homogeneous split (homo).
  • Heterogeneous Split: In heterogeneous split, we divide the dataset by subjects. The test set consists of all data from subjects 1, 2, and 9, while the training set is composed of data from the remaining 7 subjects. This design simulates a real-world scenario where the model encounters previously unseen subjects under various illumination conditions, evaluating the model’s generalization ability across different identities and lighting environments. By keeping the test subjects completely separate from the training subjects, we can ensure that the model can perform well on diverse data and avoid overfitting to specific subjects.
  • Homogeneous Split: Unlike the heterogeneous split, the homogeneous split maintains identity consistency under different illumination conditions. For each subject, we divide their sequences into three daytime/low-light sequences for training and one for validation. This design evaluates the model’s ability to handle variations of the same identity under different lighting conditions. By keeping the same subjects in both the training and validation sets, we can assess the model’s adaptability to different illuminations for known identities. Furthermore, this homogeneous split protocol validates the effectiveness of the compared methods under simpler conditions (where data similarity is high). By contrasting it with the heterogeneous split, we demonstrate that data diversity poses a greater challenge for different methods. Through comparative experimental results, we further illustrate the effectiveness of our proposed method.
5.2 Main Results on Proprietary Dataset
Figure 7 presents the normalized confusion matrix for our proposed method testing in the Heter split. The diagonal elements represent correct classifications, while off-diagonal entries indicate misclassifications. Our model demonstrates strong discriminative capabilities, particularly in wavings and numbers, achieving classification accuracies above 90%. Notable misclassifications occur primarily in freestyle motions, likely attributed to high inter-personal variability in movement execution and the test set containing completely unseen movement patterns that differ significantly from training data distributions. Other confusion occurs mainly between visually similar activities like Push forward (class 15) and Wave up (class 12) or Raise arms (class 14). All experiments are conducted on a single A100 GPU using the AdamW optimizer with a learning rate of $1\times 10^{-4}$, trained for 60 epochs with early stopping to prevent overfitting.
@@ 5.2 Main Results on Proprietary Dataset
5.3 Comparative Methods and Results
@@
  • STLLer [65]: A spiking neural network (SNN) designed for event-based action recognition, utilizing spatiotemporally aware life-long learning. This method is noted for its ability to learn continuously, adapting to new actions over time. This biologically-inspired approach achieves state-of-the-art performance on the DVS128 dataset.
@@
  • EventEgo3D [16]: The first framework addressing egocentric 3D perception tasks with event cameras, combining LNES event representation with efficient BlazeBlocks for spatial-temporal feature extraction, marking it as a pioneer in egocentric event understanding. Given its strong alignment with our task requirements, we select it as our baseline method.
@@
Table 2: Cross-Method Performance Comparison (Accuracy %).
Method Heter Split Homo Split Params (M)
Mixed Day/Night Mixed Day/Night
@@
* EventEgo3D: Originally designed for 3D keypoint estimation. We adapt its feature extractor with our classification head.
Table 2 presents the performance comparison across different methods. Our method achieves the best performance in both Heterogeneous and Homogeneous splits. For the Heterogeneous Split, our method achieves 62.7% accuracy on the Mixed-condition test, significantly outperforming other methods such as STLLer (30.2%), Assembly (59.6%), EventEgo3D (36.4%), and EventPoint (30.0%). Consistent normal light/ low light performance (62.8% day vs 62.5% night) demonstrates illumination robustness. All methods except S-tllr achieve an accuracy higher than 97% in the homo split validating experimental consistency. The parameter count of our model is 7.01M, which is relatively small, indicating that our model is efficient and effective for event-based egocentric action gesture recognition. In conclusion, our method outperforms existing approaches in action recognition under varying conditions while maintaining a relatively small number of parameters, demonstrating its robustness and effectiveness.
5.4 Component Ablation Analysis
As shown in Table 3, our ablation study reveals the effectiveness of different modules. The baseline model achieves 36.4% accuracy on the Heterogeneous Mixed testing set. Introducing the BTSM boosts the accuracy to 58.1%, demonstrating the importance of temporal information in action recognition. Adding the Visual Mamba module alone improves the accuracy to 45.3%, indicating the value of spatial feature extraction. The full model combining TSM and VMamba achieves the highest accuracy of 62.7%, showing that both modules contribute significantly to performance enhancement.
Table 3: Module Contribution in Heter Mixed Testing (%).
@@ 5.4 Component Ablation Analysis
5.5 Quantitatively Result on DVS128 Gesture
To validate the generalization capability of our proposed architecture beyond our proprietary dataset, we conduct extensive experiments on the widely recognized event-based gesture recognition benchmark, DVS128 Gesture [1]. This dataset consists of 1,342 instances of 11 distinct hand gestures captured using dynamic vision sensors, presenting unique challenges in sparse temporal feature learning. As shown in Table 4, our approach achieves a remarkable accuracy of 97.0% on this benchmark, demonstrating its strong generalization capability. This result matches the accuracy of 97.0% obtained through reimplementation of the open source state-of-the-art approach S-tllr [65] without access to its pre-trained models, despite their reported accuracy of 97.7%. For S-tllr, we follow the settings described in its official implementation. For all other methods, experiments are conducted on a single A100 GPU using the AdamW optimizer with a learning rate of $1\times 10^{-4}$, trained for 60 epochs with early stopping to prevent overfitting.
Table 4: Performance comparison on DVS128 Gesture benchmark. Our method achieves state-of-the-art accuracy without pretrained models, outperforming the reimplemented state-of-the-art S-tllr and other methods. Results highlight robustness to temporal sparsity and gesture variability. * denotes our reimplementation.

### v2

@@ III-A Head-Mounted Capture System (HMCS)
Figure 3: Action duration distributions across subjects under normal light (left) and low light (right) conditions. Each vertical bar represents the cumulative duration for a specific action, with colored segments indicating subject-wise contributions.
Our egocentric gesture acquisition platform (illustrated in Figure 1(a)) employs a custom-designed head-mounted capture system (HMCS) comprising a lightweight bicycle helmet modified with 3D-printed mounts to rigidly affix two Prophesee EVK4 event cameras and an Intel RealSense D435 depth camera. Temporal synchronization between event streams and RGB-D frames is achieved through hardware triggering, where the Intel RealSense D435 emits a synchronization signal to align data acquisition across all sensors. The event cameras utilize 4 mm focal length lenses (MV-LD-4-4M-G) providing a 94° horizontal field of view (FOV) with a native resolution of $1280\times 720$ pixels. Each pixel has a size of $4.86\mu m\times 4.86\mu m$ with temporal resolution up to 1 MHz, capturing brightness changes through asynchronous events. For normal light conditions, we employed the default EVK4 threshold configuration (0% bias_diff_off and 0% bias_diff_on). In low-light scenarios, sensitivity was enhanced by configuring 20% bias_diff_off and 20% bias_diff_on parameters to increase event generation rate. The Intel RealSense D435 RGB-D camera operates at 30 fps with synchronized $1280\times 720$ resolution for depth streams, which may be used for further research. The system connects via USB 3.0 to a host computer for power delivery and real-time data streaming, enabling uninterrupted recording sessions. The compact design and flexible HMCS allow users to freely move their heads and perform hand movements while ensuring precise temporal alignment.
III-B Data Collection Protocol
The dataset consists of data from 10 participants, encompassing a diverse range of body types and genders. Using the HMCS (Figure 1(a)), each participant performes 38 categorized gestures (Figure 1(c)) under both normal and low-light conditions (Figure 1(b)), with at least four repetitions per gesture-condition combination, ensuring a robust and varied dataset. The gestures are categorized into three main groups:
  • Single-Hand Gestures (Right Hand): 11 classes including Wave to the right, Wave to the left, Right thumb points to the right, Right index finger points to the left, Tap, Spread, Shrink, Swipe, Draw a circle, Draw an X, and NO (Shake finger).
  • Two-Hand Gestures (No Occlusion): 21 classes including Wave down, Wave up, Wave normally (arms close and move away), Raise arms, Push forward, OK, Flip palms, Grab, Make a heart, Camera gesture, Thumbs up, Show number 1-9, and Make a fist.
  • Two-Hand Gestures (Mutual Occlusion): 6 classes emphasizing challenging hand-hand interactions including Interlock fingers, Clap hands, Cross index fingers, Rotate arms, Salute with a fist, and Freestyle motion.
Figure 2 comparatively analyzes RGB and event camera outputs under varying illumination conditions, demonstrating the latter’s superior robustness to lighting variations.
III-C Dataset Composition and Statistics
The dataset provides 5,419 finely segmented event data ($\sim$193 GB) with corresponding labels based on the precise start and end times of each action. The dataset exhibits significant variability, as evidenced by two key characteristics: (1) the duration distribution of gestures varies across participants, reflecting individualized articulation speeds (Figure 3); and (2) substantial differences in normalized event rates are observed among different subjects (Figure 4). This variability in action execution patterns is crucial for training robust gesture recognition models to generalize across users and lighting. Event camera data, less lighting-sensitive than RGB, further enhances the dataset’s utility for research in challenging environments.
@@ III-C Dataset Composition and Statistics
Figure 4: Subject-wise normalized event rates showing individual adaptation variance.
@@
In summary, the established EgoEvGesture dataset offers a rich and diverse collection of egocentric gesture data, captured under varying lighting conditions, making it a valuable resource for advancing research in gesture recognition using event cameras.
IV Methodology
IV-A Architecture Overview: Dual-Motion Disentanglement Framework
@@
To address the critical challenge of ego-hand motion entanglement in event streams while capturing holistic temporal patterns, we propose a temporal pyramid architecture with three progressive processing stages (Figure 5):
@@
Figure 5: Three-stage processing pipeline: (1) LNES converts raw events to normalized surfaces preserving recent temporal signatures; (2) Blaze and VMamba blocks extract multi-scale features; (3) BTSM aligns temporal context.
@@
We first convert raw events into Locally Normalized Event Surfaces (LNES) that focus on the most recent temporal signature to suppress historical ego-motion artifacts while maintaining real-time performance. We then adopt Blaze & VMamba blocks to extract spatiotemporal features. Our Blaze encoder-decoder, adapted from [13] with proven efficiency in egocentric human perception [15], processes each temporal bin. The VMamba block addresses sparse event processing and long-term temporal modeling through discrete state space equations (detailed in Sec. IV-B). Finally, we employ the Bins-Temporal Shift Module (BTSM) to resolve temporal semantic fragmentation caused by segmented processing and perform temporal alignment (detailed in Sec. IV-C).
IV-B VMamba Block
@@
Despite the ability to adapt to high dynamic ranges, event-based gesture recognition faces challenges due to the sparsity of events, which makes it difficult for traditional methods to effectively focus on key areas. Meanwhile, the very high event resolution, while enhancing the ability to capture fine details, also makes long-term temporal understanding more challenging. To resolve these, we introduce the VMamba block, which bridges the modality gap via adaptive computation allocation — dynamically prioritizing active temporal regions while preserving temporal fidelity. Specifically, the VMamba block employs density-aware gating to route computational resources toward high-event-density intervals and suppress inactive periods, ensuring efficient feature extraction without sacrificing microsecond-level temporal granularity.
@@
IV-C Bins-Temporal Shift Module (BTSM)
@@
IV-C1 Dual-Phase Temporal Fusion
The proposed temporal fusion comprises two complementary phases sharing identical feature transformation mechanics but operating along different axes:
$\mathcal{T}(X,d)=\text{concat}\left(\begin{aligned} &\text{roll}(X_{:,:,:,0:\frac{F}{4}},\text{shift}=-1,\text{dim}=d),\\ &\text{roll}(X_{:,:,:,\frac{F}{4}:\frac{F}{2}},\text{shift}=+1,\text{dim}=d),\\ &X_{:,:,:,\frac{F}{2}:,}\end{aligned}\right)$ (1)
  • Intra-Bin Phase ($d=2$): Processes temporal relationships within bins
    $X_{intra}=\mathcal{T}(X,2)\in\mathbb{R}^{Bt\times T\times Bn\times F},$ (2)
  • Inter-Bin Phase ($d=1$): Captures cross-bin dependencies
    $X_{final}=\mathcal{T}(X_{intra},1)\in\mathbb{R}^{Bt\times T\times Bn\times F},$ (3)
Unified components:
  • $\text{roll}(\cdot)$: Cyclic shift along specified axis (negative=left, positive=right).
  • Channel split: First $F/4$ left-shifted, next $F/4$ right-shifted, last $F/2$ static.
  • Dimensions: $Bt$=batch size, $T$=bin count, $Bn$=frames/bin, $F$=features.
IV-C2 Mechanism Analysis
@@ IV-D Loss Function
V Experiments
V-A Dataset Configuration and Evaluation Protocol
@@
V-A1 Illumination Robustness Validation
We validate illumination robustness through three test scenarios:
  • Normal light: Baseline performance evaluation.
  • Low light: Assess how the model handles low light conditions.
  • Mixed: Real-world simulation with combined conditions.
V-A2 Training-Test Split Protocols
Two evaluation paradigms implemented:
  • Heterogeneous Split (Generalization):
    • Test set: Subjects 1,2,9 (unseen identities).
    • Trains on remaining 7 subjects.
    • Evaluates cross-subject robustness and the model’s generalization ability across different identities and lighting environments.
  • Homogeneous Split (Adaptability):
    • 3:1 split per subject’s sequences.
    • Maintains identity consistency and validates method effectiveness.
    • Tests illumination variation handling.
V-B Main Results on Proposed Dataset
Figure 6 presents the normalized confusion matrix for our proposed method testing in the Heter split. The diagonal elements represent correct classifications, whereas off-diagonal entries indicate misclassifications. Our model demonstrates strong discriminative capabilities, particularly in wavings and numbers, achieving classification accuracies above 90%. Notable misclassifications occur primarily in freestyle motions, likely attributed to high inter-personal variability in movement execution and the test set containing completely unseen movement patterns that differ significantly from training data distributions. Other confusion occurs mainly between visually similar activities like Push forward (class 15) and Wave up (class 12) or Raise arms (class 14). All experiments are conducted on a single A100 GPU using the AdamW optimizer with a learning rate of $1\times 10^{-4}$, trained for 60 epochs with early stopping to prevent overfitting.
@@ V-B Main Results on Proposed Dataset
V-C Comparative Methods and Results
@@
  • S-tllr [35]: A spiking neural network (SNN) designed for event-based action recognition. This biologically inspired approach achieves state-of-the-art performance on the DVS128 dataset.
@@
  • EventEgo3D [15]: The first framework addressing egocentric 3D perception tasks with event cameras, combining LNES event representation with efficient BlazeBlocks for spatial-temporal feature extraction. Given its strong alignment with our task requirements, we select it as our baseline method.
@@
TABLE II: Cross-method performance comparison (Accuracy %).
Method Heter Split Homo Split Params
Mixed Day/Night Mixed Day/Night (M)
@@
* EventEgo3D: Adapted from 3D keypoint estimation (original feature extractor + our classification head).
Table II presents the performance comparison across different methods. Our method achieves the best performance in both heterogeneous and homogeneous splits. For the heterogeneous Split, our method achieves 62.7% accuracy on the Mixed-condition test, significantly outperforming other methods such as STLLer (30.2%), Assembly (59.6%), EventEgo3D (36.4%), and EventPoint (30.0%). Consistent normal light/ low light performance (62.8% day vs 62.5% night) demonstrates illumination robustness. All methods except S-tllr achieve an accuracy higher than 97% in the homo split validating experimental consistency. The parameter count of our model is 7.01M, which is relatively small, indicating that our model is efficient and effective for event-based egocentric action gesture recognition. In conclusion, our method outperforms existing approaches in action recognition under varying conditions while maintaining a relatively small number of parameters, demonstrating its robustness and effectiveness.
V-D Ablation Study
As shown in Table III, our ablation study reveals the effectiveness of different modules. The baseline model achieves 36.4% accuracy on the Heterogeneous Mixed testing set. Introducing the BTSM boosts the accuracy to 58.1%, demonstrating the importance of temporal information in action recognition. Adding the Visual Mamba module alone improves the accuracy to 45.3%, indicating the value of spatial feature extraction. The full model combining BTSM and VMamba achieves the highest accuracy of 62.7%, showing that both modules contribute significantly to performance enhancement.
TABLE III: Module contribution in heterogeneous mixed testing (Accuracy%).
@@ V-D Ablation Study
V-E Quantitatively Result on DVS128 Gesture
To validate the generalization capability of our proposed architecture beyond our proprietary dataset, we conduct extensive experiments on the widely recognized event-based gesture recognition benchmark, DVS128 Gesture [16]. This dataset consists of 1,342 instances of 11 distinct hand gestures captured using dynamic vision sensors, presenting unique challenges in sparse temporal feature learning. As shown in Table IV, our approach achieves a remarkable accuracy of 97.0% on this benchmark, demonstrating its strong generalization capability. This result matches the accuracy of 97.0% obtained through reimplementation of the open source state-of-the-art approach S-tllr [35] without access to its pre-trained models, despite their reported accuracy of 97.7%. For S-tllr, we follow the settings described in its official implementation. For all other methods, experiments are conducted on a single A100 GPU using the AdamW optimizer with a learning rate of $1\times 10^{-4}$, trained for 60 epochs with early stopping to prevent overfitting.
TABLE IV: Performance comparison on the DVS128 Gesture benchmark. Our method achieves state-of-the-art accuracy without pretrained models, comparable to the reimplemented state-of-the-art S-tllr and other methods. Results highlight robustness to temporal sparsity and gesture variability.
* denotes our reimplementation.